A large number of bullets are fired in all directions with the same speed v. what is the maximum area of ground covered by the bullets?

Grade:11

4 Answers

Akash Kumar Dutta
10 years ago

pie.r^2
where r=v^2/g

Madhukar Thalore
10 years ago

 

Hi,

The maximum range is,
The area in which covered by the bullets will be a circle of radius equal to the maximum range (Rm)
Hence, the maximum area covered will be
Hope, this answer helps you.
Best wishes!!
vishal ojha
10 years ago

SOURAV MISHRA
10 years ago

the maximum range is given by Rmax. = v2/g.

so the maximum area covered on the ground is given by πR2max. = πv4/g2.